Two positive charges each of magnitude 5C are r metres apart in vaccum. If the force between them is 10N and C1=1/square root of 4pie permittivity. Express r in terms of C1.


1
Expert's answer
2021-04-01T18:33:18-0400

The equation for the force:


"F=C_1\frac{q·q}{r}=C_1\frac{q^2}{r^2},\\\space\\
10=C_1\frac{25}{r^2},\\\space\\
1=C_1\frac{2.5}{r^2},\\\space\\
r=2.5C_1."